What is 95=-1-6(p-8)
Found 2 solutions by patrickca, Alan3354:
Answer by patrickca(4)

Step 1: Simplify by distributing
95 = -1 + (-6(p-8))
95 = -1 + (-6p) + 48
Step 2: Combine like terms
95 = -6p + (-1 + 48)
95 = -6p +47
STep 3: SOlve using inverse ops (subtract 48 and divide by -6)
48 = -6p
-8 = p
Step 4: Check by substituting in the original equation

Answer by Alan3354(69443)

What is 95=-1-6(p-8)
================
It's an equation.